# The Waterbed Effect on Quasiperiodic Disturbance Observer: Avoidance of Sensitivity Tradeoff with Time Delays

### v1

@@ Appendix B Proof of Lemma 0
$|\varphi_{i}(j\omega)|$ $=\frac{1}{\gamma_{i}}\left(\alpha_{i}(0)+\sum_{n=1}^{N}\alpha_{i}(n)(e^{jnT\bar{U}_{i}\omega}+e^{-jnT\bar{U}_{i}\omega})\right)$
$=\frac{1}{\gamma_{i}}\left(\alpha_{i}(0)+2\sum_{n=1}^{N}\alpha_{i}(n)\cos(nT\bar{U}_{i}\omega)\right)$ (30)
@@ Appendix C Proof of Theorem 0
$\mathcal{C}_{\varepsilon}\coloneqq\{s\in\mathbb{C}\mid\mathrm{Re}[s]>\varepsilon>0\}$ (34a)
@@
$\delta\in\Delta\coloneqq\{\delta\in(\varepsilon,1/2]\mid R_{\varepsilon,\delta}>2(\delta+\omega_{\mathrm{b}})\}$ (34f)
@@
$\sup_{s\in\mathcal{A}_{\varepsilon,\delta}}|\Gamma(\delta+R_{\varepsilon,\delta}e^{j\theta})|\leq\frac{\omega_{\mathrm{b}}\omega_{\mathrm{c}}L\phi_{\varepsilon}}{R_{\varepsilon,\delta}-\delta-\omega_{\mathrm{b}}}$ (37)
$\leq 2\omega_{\mathrm{b}}\omega_{\mathrm{c}}L\phi_{\varepsilon}/R_{\varepsilon,\delta}=\delta-\varepsilon<1/2.$ (38)
@@
$\leq\frac{|\Gamma(s)|^{2}}{1-|\Gamma(s)|}\leq 2|\Gamma(s)|^{2}.$ (40)
@@
$\left|\int_{\mathcal{A}_{\varepsilon,\delta}}\log(1+\Gamma(s))-\Gamma(s)ds\right|\leq 2\pi R_{\varepsilon,\delta}\sup_{s\in\mathcal{A}_{\varepsilon,\delta}}|\Gamma(s)|^{2}$ (41)
@@
$\lim_{\varepsilon\to 0^{+}}\int_{0}^{\infty}\ln|S(\varepsilon+j\omega)|d\omega=\lim_{\varepsilon\to 0^{+}}\frac{1}{2}\int_{-\infty}^{\infty}\ln|S(\varepsilon+j\omega)|d\omega$
@@
We apply the dominated convergence theorem. Because
@@
$\lim_{\delta\to\varepsilon^{+}}\Phi(\delta+R_{\varepsilon,\delta}e^{j\theta})=0,\ \forall\theta\in\Theta,$ (48c)
the function $\Gamma(\delta+R_{\varepsilon,\delta}e^{j\theta})R_{\varepsilon,\delta}e^{j\theta}$ converges pointwise
@@
$={\omega_{\mathrm{b}}\omega_{\mathrm{c}}L}/{2}.$ (49)
Additionally,
@@
where the upper bound is an integrable function independent of $\delta$. Thus, the limit and integral can be interchanged as
$=\frac{\omega_{\mathrm{b}}\omega_{\mathrm{c}}L}{2}\int_{-\pi/2}^{\pi/2}d\theta=\frac{\omega_{\mathrm{b}}\omega_{\mathrm{c}}L\pi}{2}.$ (51)
@@ Appendix D Proof of Theorem 0
$=j\lim_{\varepsilon\to 1^{+}}\int_{0}^{2\pi}\log(\hat{S}(0))+a_{1}\varepsilon^{-1}e^{-j\theta}+\cdots d\theta$

### v2

@@ Appendix B Proof of Lemma 0
$|\varphi_{i}(j\omega)|$ $=\frac{1}{\gamma_{i}}\left|\alpha_{i}(0)+\sum_{n=1}^{N}\alpha_{i}(n)(e^{jnT\bar{U}_{i}\omega}+e^{-jnT\bar{U}_{i}\omega})\right|$
$=\frac{1}{\gamma_{i}}\left|\alpha_{i}(0)+2\sum_{n=1}^{N}\alpha_{i}(n)\cos(nT\bar{U}_{i}\omega)\right|$ (30)
@@ Appendix C Proof of Theorem 0
$\mathcal{C}_{\varepsilon}\coloneqq\{s\in\mathbb{C}\mid\mathrm{Re}[s]>\varepsilon\},\ \varepsilon\in(0,1/2)$ (34a)
@@
$\delta\in\Delta_{\varepsilon}\coloneqq\{\delta\in(\varepsilon,1/2]\mid R_{\varepsilon,\delta}>2(\delta+\omega_{\mathrm{b}})\}$ (34f)
@@
$\sup_{s\in\mathcal{A}_{\varepsilon,\delta}}|\Gamma(s)|$ $\leq\frac{\omega_{\mathrm{b}}\omega_{\mathrm{c}}L\phi_{\varepsilon}}{R_{\varepsilon,\delta}-\delta-\omega_{\mathrm{b}}}<\frac{2\omega_{\mathrm{b}}\omega_{\mathrm{c}}L\phi_{\varepsilon}}{R_{\varepsilon,\delta}}=\delta-\varepsilon$
$<1/2.$ (37)
@@
$\leq\frac{|\Gamma(s)|^{2}}{1-|\Gamma(s)|}<2|\Gamma(s)|^{2}.$ (39)
@@
$\left|\int_{\mathcal{A}_{\varepsilon,\delta}}\log(1+\Gamma(s))-\Gamma(s)ds\right|<2\pi R_{\varepsilon,\delta}\sup_{s\in\mathcal{A}_{\varepsilon,\delta}}|\Gamma(s)|^{2}$ (40)
@@
$\lim_{\varepsilon\to 0^{+}}\int_{0}^{\infty}\ln|S(\varepsilon+j\omega)|d\omega=\lim_{\delta\to 0^{+}}\frac{1}{2}\int_{-\infty}^{\infty}\ln|S(\delta+j\omega)|d\omega$
@@
The limit and integral can be interchanged on the basis of the dominated convergence theorem as
$=\frac{\omega_{\mathrm{b}}\omega_{\mathrm{c}}L}{2}\int_{-\pi/2}^{\pi/2}d\theta=\frac{\omega_{\mathrm{b}}\omega_{\mathrm{c}}L\pi}{2},$ (47)
which is based on
@@
$\lim_{\delta\to\varepsilon^{+}}\Phi(\delta+R_{\varepsilon,\delta}e^{j\theta})=0,\ \forall\theta\in\Theta$ (48c)
@@
$={\omega_{\mathrm{b}}\omega_{\mathrm{c}}L}/{2},$ (48d)
and
@@
where this upper bound is an integrable function independent of $\delta$.
@@ Appendix D Proof of Theorem 0
$=j\lim_{\varepsilon\to 1^{+}}\int_{0}^{2\pi}\log(\hat{S}(0))+a_{1}\varepsilon^{-1}e^{-j\Omega}+\cdots d\Omega$